OpenStudy (anonymous):

x^2-6=0 why isnt it (x+2) (x-3)=0 just need to know why

hartnn (hartnn):

because 3*3=9 and not 6

hartnn (hartnn):

x^2-9 = x^2-3^2 = (x+3)(x-3)

hartnn (hartnn):

x^2-6 = x^2-9+3 = (x+3)(x-3)+3

OpenStudy (anonymous):

typo its suppose to be (x-3) (x+2)

hartnn (hartnn):

ohh, (x-3)(x+2) = x^2 -3x+2x -6 =x^2 -x -6 and not x^2 -6

OpenStudy (anonymous):

if factored is it (x-3) (x+2) or square root +-6?

hartnn (hartnn):

roots are \(\large \pm \sqrt6\) factors are \((x+\sqrt6)(x-\sqrt6) \)

OpenStudy (anonymous):

but it is not (x-3) (x+2)

hartnn (hartnn):

exatly! its not because (x-3)(x+2) = x^2-x-6

OpenStudy (anonymous):

ooh i see

OpenStudy (anonymous):

(x-3) (x-2)= x^2-2x-3x+6=x^2-5x+6...so u got an extra 5x

hartnn (hartnn):

ask if still any doubts.

OpenStudy (anonymous):

so when ever it is not a perfect square, i have to to take the square root of both side?

OpenStudy (anonymous):

You have to take the square root regardless.

hartnn (hartnn):

whenever ? when its in the form of x^2 - a^2 = 0

hartnn (hartnn):

then shift a^2 to other side and take square root.

OpenStudy (anonymous):

k thanks

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

bcs u hv 1 exttra x :)

OpenStudy (anonymous):

:D